OpenStudy (anonymous):

Help Please? Nakim simplyfied (equation will be in comments) and got -10xSqrT(2x). Part |: Using complete sentences explain what nakim did wrong. Part ||: Show all your work to simpify the expression.

OpenStudy (anonymous):

\[3\sqrt{2x}+x \sqrt{8x}-5\sqrt{18x}\]

OpenStudy (reemii):

is it \(-10x \sqrt{2x}\) that is in the statement?

OpenStudy (anonymous):

Yes

OpenStudy (jdoe0001):

so, what did YOU get?

OpenStudy (anonymous):

I don't know how but i somehow got \[-2\sqrt{-16x}+x \sqrt{8x}\]

OpenStudy (anonymous):

I came on here hoping someone could explain this to me. Because to me, mine looks completely wrong.

OpenStudy (jdoe0001):

hmm, let's simplify 2 numbers there, 18 and 8 so, how else can we really rewrite 18 and 8?